OpenStudy (anonymous):

If the parallel sides of a trapezoid are contained by the lines and , find the equation of the line that contains the midsegment. Show equations and all work that leads to your answer.

hero (hero):

Please post the picture associated with the question.

OpenStudy (anonymous):

There is not a picture.

OpenStudy (anonymous):

One second, the two base are y = - 1/4 x + 5 & y = -1/4 x - 1

hero (hero):

Okay, that helps

hero (hero):

Well, the bottom base and the top base are both equidistant from the mid-segment and the slopes are all the same so it should be a matter of simply figuring out the b value by evaluating \[\frac{5 - 1}{2}\]

OpenStudy (anonymous):

okay so, 5-1 divided by 2 = 2

hero (hero):

Yes, so what is the equation of the mid-line?

OpenStudy (anonymous):

so it would be y = (-1/4)x + 2 ?

hero (hero):

Precisely